Operations Report – January 7th

Crew 202 Operations Report 07-JAN-2019

SOL: 9

Name of person filing report: Kasey Hilton

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator (hours run): 19hr 37min; Turned on last night (06Jan2019) at 13:24; Turned off this morning (07Jan2019) at 9:01; Turned on tonight (07Jan2019) at 16:26

Solar SOC – Turned on (06Jan2019) 72%; Turned off (07Jan2019) 100%; Turned on (07Jan2019) 79%

Diesel Reading – 70%

Propane Reading – 41%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 60%; 320 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 65%; 200 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 19:20 to refill tank

Water Meter: 01398787

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 66.2

Beginning charge: 100%

Ending charge: 57%

Currently charging: Yes

Opportunity rover used: No

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 64.5

Beginning charge: 100%

Ending charge: 69%

Currently charging: Yes

Notes on rovers: Brake fluid in Opportunity was replaced but the brakes are still not functioning properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: All ATVs were started to prevent the batteries from freezing

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Nothing to report

Summary of internet: Internet is not refreshing correctly; SystemControlCenter.com will show that we have no data left at the beginning of the day before anyone has turned electronics on or will show that we have 1GB but will be very slow

Summary of suits and radios: Battery in suit #9 was replaced because the battery had a voltage less than 12V

Summary of Hab operations: Side panel on west side of Hab is loose and waving in the wind, making a lot of noise; A bolt on the exhaust pipe on west side of Hab came loose

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Hot plate in ScienceDome starts producing a burning smell once the temperature is raised above 200C

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: We replaced 3 suit batteries in the last two days and replaced a charger last week, I think there are only two more back up batteries left and no back up chargers left in the suit supply box